### v2

# Decoding Human Activities: Analyzing Wearable Accelerometer and Gyroscope Data for Activity Recognition

Abstract
A person’s movement or relative positioning can be effectively captured by different types of sensors and corresponding sensor output can be utilized in various manipulative techniques for the classification of different human activities. This paper introduces a multi-structural architecture with residual networks and a spinal module developed with depth-wise separable CNN blocks, named FusionActNet, for activity recognition using sensor data. The sensor data is initially divided into two primary categories: static (where the human body is immobile) and dynamic (involving movement of the human body). Two residual networks are trained using these superclass data to effectively classify both static and dynamic activities. A spinal module consisting of depth-wise separable CNN blocks, enhanced with a guidance module, is incorporated for selecting the final prediction. Sensor data moves through the collection of these networks, which include two pre-trined ResNets, and a spinal module. After receiving prediction vectors from the static and dynamic models for the specific data, these are combined to create a concatenated vector. In this training process, the guidance module learns to generate a weight that emphasizes any of the two prediction vectors that were previously concatenated based on input data and their characteristics. The proposed model is evaluated using two publicly accessible datasets: UCI HAR and Motion-Sense. Therein, it successfully handled the highly confusing cases of data overlap. Therefore, the proposed approach achieves a state-of-the-art accuracy of 97.35% and 95.35% in the UCI HAR and Motion-Sense datasets respectively.
Index Terms— Human Activity Recognition, Deep Learning, Sensor Signal Processing, Activity Grouping, Inertial Sensors
1 Introduction
Activity recognition using wearable sensors has been a trending topic of research for its widespread applicability in diverse sectors ranging from health care services to military applications[1]. Modern mobile devices provide abundant sensor data, which are valuable for applications like activity recognition. Various types of sensor data along with image and video data have been employed for recognizing human activity[2]. In this paper, the time series wearable sensor data (e.g. 3-axis accelerometer and gyroscope) are mainly focused on, as they are pretty easy to obtain and can be used to recognize human activity from distant data which are very small in volume and easy to share through internet.
In numerous literature, support vector machine (SVM), and kth nearest neighbor (KNN) classifiers are the most popular ML-based training method for human activity recognition. Furthermore, multiple-stage training through LSTM-CNN-LSTM, ANN [3], CNN-LSTM [4], and several deep learning frameworks [5], [6] are also suggested by some literature. However, augmented-signal features and a hierarchical recognizer are combined to get a highly functional trained structure reported in [7]. In 1-D time domain data from accelerometers and gyroscopes, achieving high training accuracy is challenging with raw data. Shallow networks struggle to extract meaningful patterns, especially for activities like walking, lying, and sitting, leading to irregular patterns. Feature extraction by shallow networks often fails to provide the desired predictability. In the context of accelerometer and gyroscope data, a consistent pattern is discernible, wherein data characteristics exhibit similarities across distinct activities. For instance, during stationary states of the body such as sitting, standing, or lying down, sensor data often manifest substantial overlap. Similarly, activities involving bodily motion, such as walking upstairs, walking downstairs, or walking, also exhibit analogous data patterns. Consequently, the accurate prediction of activities presents a formidable challenge for a single network. To tackle the problem, in this work, the activities are categorized into two broad superclasses: Static (activities where the human body is stationary, e.g. sitting, laying, etc.) and Dynamic (activities where the human body is in motion, e.g. walking, jogging, etc.). This classification into super-classes is deemed essential as it allows for the creation of a distinct and effective differentiation among the various classes of activities. In this paper, a multi-structural two-stage approach that combines two residual networks with depth-wise separable CNN blocks is proposed to improve the overall performance. The key contributions of the paper are-
Introduction of a novel multi-structural two-stage network architecture, FusionActNet which leverages the strengths of residual network and dept-wise separable CNN along with a guidance module. This architecture is customized to excel in recognizing both static and dynamic activities, addressing their distinctive features effectively.
The proposed approach achieves a remarkable 97.35% accuracy by incorporating class-specific pre-trained networks for static and dynamic activities and then combining them through a spinal module consisting of depth-wise separable CNN blocks.
The remaining sections of the paper are organized in the following order. In Section 2, a brief description of the datasets used in this work, is presented. In Section 3, the problems addressed in this research work and the detailed architecture of the proposed neural network are presented. Section 4 depicts the experiment and result whereas section 5 comprises the conclusion of the paper.
2 Dataset Description
In this work, a publicly accessible dataset named UCI HAR is used [8]. It involved 30 volunteers aged 19 to 48, wearing a Samsung Galaxy S II smartphone on their waist while performing six activities which are as follows: Walking, Walking upstairs, Walking downstairs, Sitting, Standing, and Lying. Among these, Sitting, Standing, and Lying have been considered as Static activities, and Walking, Walking upstairs, and Walking downstairs have been considered as Dynamic activities for this work. The smartphone’s accelerometers and gyroscopes recorded 3-axial linear acceleration and angular velocity at 50Hz. The dataset was split into 70% training and 30% testing data. Pre-processing included noise filtering, fixed-width sliding windowing, and separating gravitational and body motion components. The dataset comprised 7352 training observations from 30 subjects and 2947 test observations for model evaluation.
Another dataset used in this work is the Motion-Sense dataset, which consists of 15 trials involving 24 individuals using an iPhone 6s placed in their trouser pockets to record accelerometer and gyroscope data [9]. It includes 12 features, such as attitude, gravity, rotation rate, and user acceleration, and covers activities like Sitting, Standing, Walking, Walking Downstairs, Walking Upstairs, and Jogging. This dataset contains 2 static classes and 4 dynamic classes.
3 METHODOLOGY
Within the domain of analyzing mixed-class activities characterized by significant overlap in data patterns, the application of deep neural networks for feature extraction confronts challenges in achieving robust generalization across diverse datasets. The challenge becomes more enhanced when gyroscope and accelerometer data are used for the identification of human activities, specifically closely similar activities, such as walking, walking upstairs, or walking downstairs. In this case, activities display significant similarities in their respective data patterns. As a result, the features retrieved from these signals become strongly correlated, which poses a considerable challenge while appropriately categorizing these activities using a single deep learning network. To address this issue, the concept of categorizing human activities into two superclasses: Static and Dynamic, is introduced here. At first, these two superclass data are trained by two separate models, allowing each model to specialize in a specific type of data. Instead of relying on a single network to be trained, this approach leverages two specialized models, making it more effective in capturing subtle patterns within the data from these two superclasses. During the second training step, these two pre-trained models provide predictions for a specific input data, which are then combined using guidance-based weighted concatenation to ensure proper superclass assignment and improve prediction accuracy.
3.1 Problem Formulation
In mathematical terms, a given data set of accelerometer and gyroscope is denoted by where is the number of samples, denotes the inertial signal data and represents the activity label. As mentioned before, the dataset is divided into two activity superclasses - static and dynamic data, each comprising the samples of all the static and dynamic activity classes, respectively. Let us denote static data as and dynamic data as , where . At the initial training stage, a static model is trained using and a dynamic model is trained using . After the first stage training process, two expert models are obtained that are specialized at classifying static and dynamic activity data respectively.
At the next stage, a data sample , which might be either from a static or dynamic superclass, is passed through both trained models and the generated prediction vectors can be written as,
Let symbolize the prediction vector of the static model and represent the prediction vector of the dynamic model for the signal in this context. It is important to note that due to the distinct attributes of the models and , usually only one of the prediction vectors or would be relevant for the analytical objectives. However, both prediction vectors are included in the analysis at this point, as the final assignment of the sample to a superclass is still unknown. In order to get the desired output, a guidance factor is needed, which is obtained by processing the input with the spinal module and controller block as shown in Fig. 1. Then is combined with previously obtained and to calculate the output activity label probabilities formulated as,
with a loss function
Here is the true activity class. The goal is to maximize the model’s accuracy and dependability by minimizing the loss function mentioned in eqn. 3.
3.2 Network Architecture
In this study, an effective deep neural network architecture has been proposed, that utilizes the weighted concatenation of outputs from two expert models and , to recognize static and dynamic human activities respectively. Because of this weighted fusion of outputs from different expert models to recognize human activities from different sensor data, this network is named FusionActNet. The proposed model leverages the potential of 1D residual blocks as well as depth-wise separable building blocks, offering a dynamic and static pathway for information processing.
3.2.1 Static Pathway
The static component of the FusionActNet is used for identifying the static human activities (e.g. sitting, laying, and standing) that comprise a series of residual blocks, as shown in Fig. 2, each of which consists of a convolutional layer, batch normalization, and rectified linear unit (ReLU) activation functions as well as a skip connection. Each residual block is represented by where and refers to the input and output dimensions of 1D convolution layers and refers to the kernel size of the 1D max-pooling layer. The residual block begins with a convolutional layer with a 3x3 kernel and output channels , followed by batch normalization and ReLU activation. Subsequently, another 3x3 convolutional layer with output channels is applied, followed by batch normalization. An identity convolutional layer with a 1x1 kernel is used to match dimensions for residual connections. A max-pooling layer with a down-sampling factor is applied next. The final layer of the static pathway includes average pooling and a fully connected layer to reduce the output to output features, where, indicates the number of static activity labels.
3.2.2 Dynamic Pathway
The dynamic pathway is structurally identical to the static pathway but is expected to capture dynamic patterns or temporal information in the input dynamic data. It follows the same residual block structure but is trained with dynamic human activity data such as walking, walking upstairs, and walking downstairs, having an equal number of output features, to the number of dynamic activity labels in the dataset.
3.2.3 Spine Module
The neural module at the center of Fig. 2 is called the Spine module, followed by a series of depthwise separable convolution blocks, DwSep, that consists of a 1D convolution layer, residual skip connection, the batch normalization, and ReLU activation. A 3x3 depth-wise convolution is applied for model size reduction and improving efficiency, followed by batch normalization and ReLU activation. A 1x1 point-wise convolution is used for dimensionality reduction, followed by batch normalization.
3.2.4 Controller Module,
After traversing the spine module, the input proceeds into the Controller Module, a pivotal component in our architecture. In the Controller Module, the guidance block encompasses three key operations: average pooling, a linear layer, and a sigmoid block. Following the execution of these operations, a guidance factor denoted as , is generated. The dimensionality of is (1,1) as the context factor is prioritizing the output labels of a specific superclass, either or . In other terms, the guidance factor can function as a classifier to discern whether the input data pertains to a static or dynamic context.
Upon passage through class-specific pre-trained static and dynamic models, the resultant outputs and are consolidated through concatenation. Subsequently, following the generation of the guidance factor , a weighted concatenation is executed on the previously merged output of and . This strategic approach is employed to accord priority to the output of the pre-trained expert model. In simple terms, during the cases of static activity classes, the guidance factor will assign higher priority to the output of . Similarly, the outputs of will be multiplied by larger weights in the cases of dynamic activity classes, whereas the prediction of the other model will be assigned a very small weight. The deliberate fusion of these outputs enhances the model’s adaptability and robustness, endowing it with the capability to discern and classify activities characterized by diverse and nuanced features. This fusion process, driven by learned weights, contributes to the model’s efficacy in recognizing and categorizing activities with varying characteristics, thus fortifying its utility in dynamic and real-world scenarios.
4 RESULTS AND DISCUSSIONS
Method | Accuracy | Precision | Recall | F1-score |
SSAL [10] | 0.8124 | 0.8166 | 0.8125 | 0.8145 |
MC-HF-SVM [11] | 0.8935 | 0.8993 | 0.8968 | 0.8980 |
ResNet [12] | 0.9073 | 0.9095 | 0.8998 | 0.9046 |
Res-BiLSTM [13] | 0.9160 | 0.915 | - | - |
Bi-LSTM [14] | 0.9270 | 0.9270 | - | - |
CNN-LSTM Self-Att. [15] | 0.9310 | 0.9090 | - | - |
CNN [16] | 0.9271 | 0.9321 | 0.9321 | 0.9293 |
HAR-CT: CNN [17] | 0.9406 | 0.9358 | 0.9359 | 0.9359 |
iSPLInception [15] | 0.951 | 0.95 | - | - |
GRU+Attention [18] | 0.96 | 0.958 | - | - |
CNN-DCT [19] | 0.971 | - | - | - |
FusionActNet | ||||
(Proposed) | 0.9735 | 0.97 | 0.97 | 0.97 |
Method | Accuracy | Precision | Recall | F1-score |
MFCC + SVM [20] | 0.8935 | 0.8993 | 0.8968 | 0.8980 |
SVM + PSO [21] | 0.8124 | 0.8166 | 0.8125 | 0.8145 |
GA + ACO [21] | 0.9073 | 0.9095 | 0.8998 | 0.9046 |
DT + BGWO [21] | 0.9271 | 0.9321 | 0.9321 | 0.9293 |
Self-supervised TPN [22] | 0.8901 | 0.8901 | 0.8899 | 0.8899 |
FusionActNet | ||||
(Proposed) | 0.9535 | 0.95 | 0.95 | 0.95 |
In this section, the performance of various methods is evaluated using the UCI-HAR and MotionSense datasets with four conventional evaluation criteria, namely accuracy, precision, recall, and F1-score. For the purpose of performance comparison, some state-of-the-art methods are considered and results are presented along with their corresponding evaluation metrics in Table 1 for the UCI-HAR dataset and Table 2 for the MotionSense dataset. The trend seems to favor multilayer perceptron or neural network-based approaches as an obvious choice for model prediction of activity recognition. To be precise, CNN-DCT with a respective accuracy of 97.1% called the closest shot with our proposed FusionActNet with an accuracy of 97.35%. In order to present the performance in each class, the confusion matrixes of corresponding datasets are shown in Table 3. From Table 3 it is observed in a confusion matrix that activity class ’walking’ is misclassified as ’walking upstairs’ and ’walking downstairs’ which are very logical. The misclassified cases are investigated closely and it is found that in these cases, the nature of the data is quite similar. Another point of limitation is found in the ’sitting’ and ’standing’ classes where the data are very similar due to the highly static and overlapping nature.
WA | WU | WD | SI | ST | LA | |
WA | 0.95 | 0 | 0.05 | 0 | 0 | 0 |
WU | 0.01 | 0.98 | 0.01 | 0 | 0 | 0 |
WD | 0 | 0 | 0 | 0.93 | 0.07 | 0 |
SI | 0 | 0 | 0 | 0.94 | 0.06 | 0 |
ST | 0 | 0 | 0 | 0.02 | 0.98 | 0 |
LA | 0 | 0 | 0 | 0 | 0 | 1.00 |
SI | ST | WD | WU | JG | WA | |
SI | 1.00 | 0 | 0 | 0 | 0 | 0 |
ST | 0 | 1.00 | 0 | 0 | 0 | 0 |
WD | 0 | 0 | 0.82 | 0.07 | 0.07 | 0.04 |
WU | 0 | 0 | 0.05 | 0.86 | 0.07 | 0.02 |
JG | 0 | 0 | 0.02 | 0.02 | 0.95 | 0.01 |
WA | 0 | 0 | 0.03 | 0.02 | 0.03 | 0.92 |
5 Conclusion
This paper presents FusionActNet, a multi-structural architecture employing residual networks and a spinal module with depth-wise separable CNN blocks for activity recognition using sensor data. For effectively classifying static and dynamic activities two distinct models are employed to capture the underlying pattern of closely related activities. Additionally, a guidance module is included for selecting the final prediction vectors, which are generated from the pre-trained models. The proposed FusionActNet achieves state-of-the-art accuracy on both UCI-HAR and Motion-Sense datasets, demonstrating robustness in handling data overlap scenarios. This approach showcases promising advancements in activity recognition techniques, with potential applications in various domains requiring precise human activity classification.